write a system of two equations in two unknowns. Ziggy's famous yougurt blends regular yogurt that is 3% fat with its no fat yogurt to obtain low fat yogurt that is 1% fat. How may pounds of regular yogurt and how many pounds of no fat yogurt should be mixed to obtain 60 pounds of low fat yogurt?
Answer by stanbon(57407)

write a system of two equations in two unknowns. Ziggy's famous yougurt blends regular yogurt that is 3% fat with its no fat yogurt to obtain low fat yogurt that is 1% fat. How may pounds of regular yogurt and how many pounds of no fat yogurt should be mixed to obtain 60 pounds of low fat yogurt?
---------------------
Equations:
Quantity Eq.: x + y = 60
Fat Eq.....:0.03x+0*y = 0.01(60)
=====================================
Multiply thru the "Fat" equation by 100 to get:
x + y = 60
3x = 60
===============
x = 20 lbs (amt. of 3% yogurt needed in the mixture)
y = 60-20 = 40 lbs (amt. of 0% yogurt needed in the mixture)
===============================================================
Cheers,
Stan H.